# Hamiltonian and Liouvillian learning in weakly-dissipative quantum many-body systems

### v1

@@ IV Conclusion and Outlook
In this study, we have devised a method for learning Hamiltonian and Liouvillian in the analog quantum simulation of many-body systems. Our work applies to a scenario where one has direct access to the quantum device, however, in the literature, other scenarios have been considered, where one does not have direct access to the quantum device, see, e.g., Refs. [25, 52]. Our method is applicable in a regime of experimental relevance where the dissipation is weak compared to the coherent evolution. Our protocol is based on quench experiments, where initial product states evolve under coherent and dissipative dynamics, and the resulting state is measured in a product basis. Hamiltonian and Liouvillian learning can be understood as a sample efficient process tomography of quantum simulators. The learning method begins with an ansatz for the operator structure of the Hamiltonian and Liouvillian. The quality of this ansatz can be monitored by measuring the learning error. Our strategy encompasses the reparametrization of the ansatz. This typically allows data recycling from previous measurement runs, but requires additional classical post-processing. This approach enables us to identify step by step first the dominant operator content of the Hamiltonian and Liouvillian, and successively sub-dominant terms within a limited measurement budget. A distinctive feature of our approach is that we can learn the Hamiltonian without the necessity to learn the entire Liouvillian, thus reducing the number of parameters to be learned. However, we demonstrated that additional constraints can be employed to learn the entire Liouvillian and to ascertain the overall scale of the Hamiltonian. Furthermore, these additional constraints can facilitate the learning of the Hamiltonian even when there are conserved quantities as operators commuting with the system Hamiltonian. While the focus of the present paper has been on spin models, the central ideas of Hamiltonian and Liouvillian learning also carry over to Bose and Fermi Hubbard models.
@@
Extensions of the present work should consider scenarios where the experimental Hamiltonian (and Liouvillian) involves a large number of small terms, which, e.g., emerge as corrections in effective many-body spin models in a low-energy description. Such a formulation might involve a statistical description as learning of an ensemble of Hamiltonians. Along similar lines, Hamiltonian learning might also account for slow drifts of experimental Hamiltonians and Liouvillians. Considering alternative viewpoints, Hamiltonian and Liouvillian learning can also be phrased in the language of Bayesian inference, similar to Ref. [30], establishing an interesting link between techniques of parameter estimation in multi-parameter quantum metrology, and optimal sensing with finite measurement budgets. Finally, exploring alternative routes should include shadow-tomography tomography [53] to estimate expectation values of many few-body observables, relevant in cases where the Hamiltonian consists of many non-commuting terms, and using (short-range) entangled states as inputs, which may be easily prepared in experiments.
@@ Appendix A Additional constraints
As we have already pointed out in the main text, Eq. (7) alone is not sufficient for learning the Hamiltonian, as $H$ is only determined up to an overall scale, and, possibly, conserved quantities not proportional to $H$. Moreover, it may happen that the dissipation rates are not uniquely determined, even in case the ansatz for the dissipation is correct. These ambiguities result from invariances of the conditions in Eq. (7), some of which we have already discussed in the main text. In the following, we show that by including additional constraints Eq. (7) can become sufficient for learning $H$, including its overall scale, and to learn the Liouvillian.
As explained in the main text, the idea is to impose additional constraints to single out the vector $\bm{c}^{H}$ of $H$ (or the vector $\bm{d}^{\mathcal{L}}$ of dissipation rates of $\mathcal{L}_{D}$) as the unique solution to Eq. (9). To this end, we consider the equation of motion of a general observable, $O$, not commuting with $H$, which is given by the Ehrenfest theorem, i.e.,
$\frac{\mathrm{d}}{\mathrm{d}t}\expectationvalue*{O}=\expectationvalue{-i[O,H]}\\ +\frac{1}{2}\sum_{k}\gamma_{k}\langle l_{k}^{\dagger}[O,l_{k}]+[l_{k}^{\dagger},O]l_{k}\rangle.$ (39)
Instead of choosing $O=H$, as we did below Eq. (3) to obtain the constraints in $M(\bm{d})$, we now have to insert the ansätze $A(\bm{c})$ for $H$ and $\Gamma(\bm{d})$ for the dissipative terms. Thus, after taking the integral from $t=0$ to $t=T$, we obtain a set of linear equations for $\bm{c}$ and $\bm{d}$. Again, these can be written as a simple matrix equation;
@@
Figure 5: (a) Singular value spectrum of $M_{G}(\beta)$ defined in Eq. (51) as a function of the regularization strength $\beta$ for the model introduced in Section III.2 in the main text. Increasing $\beta$ opens a gap between the image of the parametrization matrix $G$ introduced in Appendix B and the rest of the spectrum. Within this subspace, $\lambda(\bm{c}^{\mathrm{rec}})$, defined in Eq. (52), follows the best approximation to the Hamiltonian of the system. (b) Scaling of the second-lowest singular value $\lambda_{2}$ with system size $N$, for the same model and ansätze as in Section III.1 and fixed $N_{\mathrm{runs}}=3.1\times 10^{7}$. One finds that $\lambda_{2}$ does not decrease with system size if the number of free parameters in the ansatz is kept constant, which is the case for all ansätze except $A_{5}$.
Appendix B (Re-)Parameterization
@@
As we have seen in the main text, ansatz reparamtetrization is one of the central tools used in the strategy we have devised. Therefore, we want to discuss here in more detail the different possibilities to reparametrize an ansatz. In general, there are two different strategies but we will show that both can be understood in terms of regularization of the cost function in Eq. (9).
@@
In the first reparametrization strategy the vector of parameters, $\bm{c}$, is mapped via a parametrization matrix $G$ to a new vector of parameters, $\bm{c}_{G}=G^{T}\bm{c}$. Here, the matrix $G$ is a $n\times\tilde{n}$ matrix, with $n\geq\tilde{n}$, encoding the dependencies between the $n$ parameters in $\bm{c}$, and the $\tilde{n}$ parameters in $\bm{c}_{G}$. Moreover, one requires that $G=(\bm{g}_{1},\dots,\bm{g}_{\tilde{n}})$, with orthonormal columns $\langle\bm{g}_{i}|\bm{g}_{j}\rangle=\delta_{ij}$. One can easily verify that $G^{T}G=\mathds{1}_{\tilde{n}\times\tilde{n}}$, and $GG^{T}=\sum_{i=1}^{\tilde{n}}\outerproduct{\bm{g}_{i}}{\bm{g}_{i}}$ |, which is a projector onto the support of $G^{T}$, and thus $G$ is an isometry. In particular cases $G$ can also depend on non-linear parameters, i.e.,
@@
for the example in Section III.2 (see Fig. 5a). As expected, in the limit $\beta\to 0$, the spectrum converges to the spectrum of the unparametrized constraint matrix $M(\bm{d}^{\rm rec})$. Here, the gap between the lowest singular values and the rest of the spectrum becomes very small, resulting in an unstable solution. Then, when increasing $\beta$, all singular values, whose corresponding right-singular vectors are incompatible with the regularization increase with $\beta$, opening a gap to the subspace spanned by the regularization, i.e., the image of $G$, which in the case of the ansatz in Eq. (33) is $2$-dimensional. Note, that the Hamiltonian in Eq. (28) only approximately lies in the image of $G$. Therefore, also the lowest singular values initially increase with $\beta$, until they reach a constant value, that corresponds to an exactly enforced parametrization. Along this path, we can monitor the minimum of the cost function in Eq. (50) defined via
@@
which follows the singular value that corresponds to the best approximation of the Hamiltonian. This indicates successful exclusion of the conserved quantity via the additional constraints, as discussed in Section III.2 in the main text.
Figure 6: Comparison of $\lambda_{1}/\lambda_{2}$ (experimentally accessible), and $\sin(\theta)$ (not experimentally accessible), as a function of the number of runs, $N_{\rm runs}$. The different ansätze are defined in Section III.1 in the main text. Error bars are computed via bootstrapping as described in Appendix C, using $80$ samples. The colored triangles indicate the asymptotic values in the limit $N_{\rm runs}\to\infty$. Both quantities show a similar behaviour which supports using $\lambda_{1}/\lambda_{2}$ to asses the quality or learning error of an ansatz.
Appendix C Bootstrapping
@@
In our case the $X_{1},\dots,X_{n}$ are the individual measurements in a given basis for fixed initial state and simulation time. One can also think of $X_{1},\dots,X_{n}$ to be individual measurements of a given observable for fixed initial state and simulation time, in the case where the measurements of different observables are independent. Then the quantities of interest $T(X_{1},\dots,X_{n})$ are, e.g., the learned parameters $\bm{c}^{\mathrm{rec}}$, or the ratio of singular vales $\lambda_{1}/\lambda_{2}$. The number of samples $r$ is chosen the minimum possible integer, such that the error bars do not significantly change anymore when further increasing $r$.
Appendix D Scalability
In the context of Hamiltonian learning it was shown in Ref. [35] that, under the premise that $H$ is a local many-body Hamiltonian, the experimental resources required for Hamiltonian learning scale only polynomially in the system size, and are even constant in system size in the case of a translation invariant Hamiltonian. As we are considering the limit of weakly dissipative dynamics, where the Liouvillian only gives a small perturbation to the coherent time-evolution generated by the many-body Hamiltonian, we expect a similar scaling to apply. More precisely, one expects that the experimental resources, such as, for instance, the number of measurements, required to obtain a certain accuracy, depend on the total number of parameters of the learning task. In Fig. 5b, we show the stability of the Hamiltonian reconstruction in the presence of weak dissipation, characterized by the gap of the constraint matrix $M(\bm{d}^{\mathrm{rec}})$, for the model system discussed in Section III.1 in the main text, as a function of the system size $N$, while keeping constant the total measurement budget $N_{\rm runs}$. We find that, for ansätze with a constant number of parameters, some of which are insufficient while others are sufficient, the gap remains approximately constant in system size, while for an unparamtetrized ansatz, where the number of parameters grows with $N$, the gap closes as the system size increases. This highlights the necessity of learning an efficient parametrization of the Hamiltonian when scaling Hamiltonian and Liouvillian learning to larger system sizes.
Appendix E Comparison between $\lambda_{1}/\lambda_{2}$ and $\sin(\theta)$

### v2

@@ IV Conclusion and Outlook
In this study, we have devised and compared different methods for learning the Hamiltonian and Liouvillian in the analog quantum simulation of many-body systems. Our work applies to a scenario where one has direct access to the quantum device, however, in the literature, other scenarios have been considered, where one does not have direct access to the quantum device, see, e.g., Refs. [25, 55]. Our method is applicable in a regime of experimental relevance where the dissipation is weak compared to the coherent evolution. Our protocols are based on quench experiments, where initial product states evolve under coherent and dissipative dynamics, and the resulting state is measured in a product basis. Hamiltonian and Liouvillian learning can be understood as a sample efficient process tomography of quantum simulators. The learning methods begin with an ansatz for the operator structure of the Hamiltonian and Liouvillian. The quality of this ansatz can be monitored by measuring the learning error. Our strategy encompasses the reparametrization of the ansatz. This typically allows data recycling from previous measurement runs, but requires additional classical post-processing. This approach enables us to identify step by step first the dominant operator content of the Hamiltonian and Liouvillian, and successively sub-dominant terms within a limited measurement budget. A distinctive feature of learning from generalized energy conservation is that we can learn the Hamiltonian without the necessity to learn the entire Liouvillian, thus reducing the number of parameters to be learned. However, we demonstrated that additional constraints can be employed to learn the entire Liouvillian and to ascertain the overall scale of the Hamiltonian. Furthermore, these additional constraints can facilitate the learning of the Hamiltonian even when there are conserved quantities as operators commuting with the system Hamiltonian. We demonstrate how randomized measurements [47] can be utilized to learn the Hamiltonian and Liouvillian from the generalized Ehrenfest theorem, as this requires the estimation of many different correlation functions. Moreover, this allowed us to numerically demonstrate Hamiltonian learning for large systems consisting of tens of spins. While the focus of the present paper has been on spin models, the central ideas of Hamiltonian and Liouvillian learning also carry over to Bose and Fermi Hubbard models.
@@
Extensions of the present work should consider scenarios where the experimental Hamiltonian (and Liouvillian) involves a large number of small terms, which, e.g., emerge as corrections in effective many-body spin models in a low-energy description. Such a formulation might involve a statistical description as learning of an ensemble of Hamiltonians. Along similar lines, Hamiltonian learning might also account for slow drifts of experimental Hamiltonians and Liouvillians. Considering alternative viewpoints, Hamiltonian and Liouvillian learning can also be phrased in the language of Bayesian inference, similar to Ref. [31], establishing an interesting link between techniques of parameter estimation in multi-parameter quantum metrology, and optimal sensing with finite measurement budgets. Finally, exploring alternative routes should include using (short-range) entangled states as inputs, which may be easily prepared in experiments.
@@ Appendix A Additional constraints
We consider the equation of motion of a general observable, $O$, not commuting with $H$, as given in Eq. (4a). Inserting the ansatz $A(\bm{c})$ for $H$ and $\mathcal{D}(\bm{d})$ for the Liouvillian one obtains again a simple matrix equation;
@@
Appendix B Effect of shot noise
Here, we discuss in more detail the role of the two smallest singular values of the constraint matrix. As an example we will consider Eq. (11), but similar arguments also apply to Eq. (7).
In case the ansatz for the dynamics is sufficient, the only source of error is shot noise. Therefore, with a finite measurement budget one obtains a noisy estimate, $\widetilde{M}(\bm{d})=M(\bm{d})+E(\bm{d})$, of the true constraint matrix $M(\bm{d})$, with an additive error matrix $E(\bm{d})$. One can establish the following bound on the perturbed singular value $\lambda_{1}$
$\lambda_{1}[\widetilde{M}(\bm{d}^{\mathrm{rec}})]\leq\lambda_{1}[\widetilde{M}(\bm{d}^{\gamma})]\leq\norm{E(\bm{d}^{\gamma})},$ (39)
where the first inequality holds as $\bm{d}^{\mathrm{rec}}$ is the minimum of the cost-function in Eq. (14), and the second inequality follows from Weyl’s inequality and the fact that $\lambda_{1}[M(\bm{d}^{\gamma})]=0$ for a sufficient ansatz. Therefore, in early stages of the learning procedure, one expects $\lambda_{1}$ to decrease $\sim N_{\rm runs}^{-1/2}$ as we increase the number of runs, $N_{\rm runs}$, in the experiment. As the ansatz is sufficient $\lambda_{1}[\widetilde{M}(\bm{d}^{\mathrm{rec}})]$ will be zero in the absence of shot noise. However, if an ansatz is insufficient, $\lambda_{1}$ is strictly bounded away from zero, even in the absence of shot noise. Thus, for $N_{\rm runs}$ sufficiently large, $\lambda_{1}$ will reach a plateau. Here, the dominant source of error will be systematic errors due to missing terms in our ansatz.
$|\sin(\theta)|\leq\frac{\norm{E(\bm{d}^{\mathrm{rec}})+R(\Delta\bm{d})}}{\delta},$ (40)
where $R(\Delta\bm{d})=\sum_{k}\Delta d_{k}M_{D}^{(k)}$. Here, $\delta=\lambda_{2}-\lambda_{1}^{\rm exact}$, and $\lambda_{1}^{\rm exact}=0$ for a sufficient ansatz. In case the elements of $E$ are i.i.d. Gaussian random variables one can establish average bounds on $|\sin(\theta)|$ [36]. However, we emphasize that the noise will in general be correlated due to the fact that we perform many of the required measurements in parallel. Moreover, the error matrix $E(\bm{d})$ is experimentally inaccessible. Nevertheless, as we demonstrate in Section III, the ratio $\lambda_{1}/\lambda_{2}$, which can be directly computed from experimental data, serves as a quantity to asses an ansatz.
Appendix C (Re-)Parameterization
@@ Appendix B Effect of shot noise
Figure 6: Singular value spectrum of $M_{G}(\beta)$ defined in Eq. (46) as a function of the regularization strength $\beta$ for the model introduced in Section III.3 in the main text. Increasing $\beta$ opens a gap between the image of the parametrization matrix $G$ introduced in Appendix C and the rest of the spectrum. Within this subspace, $\lambda(\bm{c}^{\mathrm{rec}})$, defined in Eq. (47), follows the best approximation to the Hamiltonian of the system (dashed line).
@@
We want to discuss here in more detail the different possibilities to reparametrize an ansatz. In general, there are two possibilities, but we will show that both can be understood in terms of regularization of the cost function in Eq. (14). Note, that the same holds for Eq. (10).
@@
In the first reparametrization the vector of parameters, $\bm{c}$, is mapped via a parametrization matrix $G$ to a new vector of parameters, $\bm{c}_{G}=G^{T}\bm{c}$. Here, the matrix $G$ is a $n\times\tilde{n}$ matrix, with $n\geq\tilde{n}$, encoding the dependencies between the $n$ parameters in $\bm{c}$, and the $\tilde{n}$ parameters in $\bm{c}_{G}$. Moreover, one requires that $G=(\bm{g}_{1},\dots,\bm{g}_{\tilde{n}})$, with orthonormal columns $\langle\bm{g}_{i}|\bm{g}_{j}\rangle=\delta_{ij}$. One can easily verify that $G^{T}G=\mathds{1}_{\tilde{n}\times\tilde{n}}$, and $GG^{T}=\sum_{i=1}^{\tilde{n}}\outerproduct{\bm{g}_{i}}{\bm{g}_{i}}$ |, which is a projector onto the support of $G^{T}$, and thus $G$ is an isometry. In particular cases $G$ can also depend on non-linear parameters, i.e.,
@@
Finally, one notices that given that $GG^{T}\bm{c}^{H}=\bm{c}^{H}$, i.e., the parametrization does not render an ansatz insufficient, the bound in Eq. (40) can only tighten. To see this, one observes that $\lambda_{2}[M_{G}(\bm{d})]\geq\lambda_{2}[M(\bm{d})]$, and $\norm{E(\bm{d})G}\leq\norm{E(\bm{d})}\norm{G}=\norm{E(\bm{d})}$ for any isometry $G$. Therefore, reducing the number of parameters in general increases the gap $\delta=\lambda_{2}$ of the constraint matrix $M(\bm{d})$.
Figure 7: Comparison of $\lambda_{1}/\lambda_{2}$ (experimentally accessible), and $|\sin(\theta)|$ (not experimentally accessible), as a function of the number of runs, $N_{\rm runs}$. The different ansätze are defined in Section III.2 in the main text. Error bars are computed via bootstrapping as described in Appendix D, using $80$ samples. The colored triangles indicate the asymptotic values in the limit $N_{\rm runs}\to\infty$. Both quantities show a similar behaviour which supports using $\lambda_{1}/\lambda_{2}$ to asses the quality or learning error of an ansatz.
@@
for the model system in Section III.3 (see Fig. 6a). As expected, in the limit $\beta\to 0$, the spectrum converges to the spectrum of the unparametrized constraint matrix $M(\bm{d}^{\rm rec})$. Here, the gap between the lowest singular values and the rest of the spectrum becomes very small, resulting in an unstable solution. Then, when increasing $\beta$, all singular values, whose corresponding right-singular vectors are incompatible with the regularization increase with $\beta$, opening a gap to the subspace spanned by the regularization, i.e., the image of $G$, which in the case of the ansatz in Eq. (28) is $2$-dimensional. Note, that the Hamiltonian in Eq. (23) only approximately lies in the image of $G$. Therefore, also the lowest singular values initially increase with $\beta$, until they reach a constant value, that corresponds to an exactly enforced parametrization. Along this path, we can monitor the minimum of the cost function in Eq. (45) defined via
@@
which follows the singular value that corresponds to the best approximation of the Hamiltonian.
Appendix D Bootstrapping
@@
In our case the $X_{1},\dots,X_{n}$ are the individual measurements in a given basis for fixed initial state and simulation time. One can also think of $X_{1},\dots,X_{n}$ to be individual measurements of a given observable for fixed initial state and simulation time, in the case where the measurements of different observables are independent. Then the quantities of interest $T(X_{1},\dots,X_{n})$ are, e.g., the learned parameters $\bm{c}^{\mathrm{rec}}$, or the ratio of singular values $\lambda_{1}/\lambda_{2}$. The number of samples $r$ is chosen the minimum possible integer, such that the error bars do not significantly change anymore when further increasing $r$.
Appendix E Comparison between $\lambda_{1}/\lambda_{2}$ and $|\sin(\theta)|$
@@
Appendix F Learning from generalized energy conservation and additional constraints
Figure 8: Hamiltonian learning in presence of collective dissipation. (a) The ratio $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}$ after excluding the conserved quantity $Q$ for different regularization parameters $\beta$. The colored triangles describe the asymptotic values at $N_{\mathrm{runs}}\to\infty$. The dashed line follows the learning procedure described in the main text. (b) Snapshots of the learned Hamiltonian parameters, $c^{\rm rec}_{j}$, corresponding to the parameters $J_{kl}$ and $B_{i}$ defined in Eq. (23). The required measurement budget is also indicated in the left panel. Black crosses correspond to the parameters of the model Hamiltonian specified below. All error bars are computed via bootstrapping as explained in Appendix D, using $10$ samples.  (13) are evaluated using $B_{z}\Delta_{T}=1/128$, i.e., using $128$ time-steps, where at each time we use only a fraction of $1/128$ number of measurements per basis, compared to the quench times $B_{z}T\in[0,0.5,1]$. Expectation values at $t=0$ are calculated exactly from the initial state. The auxiliary parameter $\xi$ for the additional constraints as introduced in Eq. (15) is set to $\xi=1000$.
In the following we compare the learning based on the generalized Ehrenfest theorem, presented in Section III.3 in the main text, to the leaning based on energy conservation. We consider the same model systems comprising the long-range Hamiltonian $H_{\mathrm{XY}}$ as well as local and collective dissipation, introduced in Section III.3 in the main text.
We will use Eq. (28) as an ansatz for the Hamiltonian. As we also want to learn the dissipative processes, we will start with an ansatz for the Liouvillian, comprising single-qubit Lindblad operators that typically appear in the context of trapped-ion experiments, i.e.,
$\mathcal{D}^{\mathrm{loc}}(\varrho)=\sum_{\eta\in\{+,-,z\}}\frac{d_{\eta}}{2}\sum_{k=1}^{N}\quantity([\sigma_{k}^{\eta}\varrho,\sigma_{k}^{\eta,\dagger}]+\mathrm{h.c.}).$ (48)
Firstly, one notices that the ansatz $A$ contains the total magnetization defined by
$Q=B\sum_{k}\sigma^{z}_{k},$ (49)
In Fig. 8a we monitor $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}$ as a function of $N_{\mathrm{runs}}$ for the ansatz $A$, and for different values of the regularization parameter $\beta$. One starts with a large value of the regularization parameter, here, $\beta=100$, which strongly imposes the parametrization of $A$, leading to reconstructed parameters without spatial disorder, see Fig. 8b for $N_{\mathrm{runs}}=2.1\times 10^{5}$ (blue data). Then further increasing $N_{\mathrm{runs}}$ reduces the size of the error bars in $\bm{c}^{\mathrm{rec}}$, as is also shown in Fig. 8b for $N_{\mathrm{runs}}=1.7\times 10^{6}$ (blue data). However, an ansatz with $\beta=100$ cannot account for disorder in the coupling terms in $H_{XY}$, which is shown by the plateau in $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}$, here, starting at around $N_{\mathrm{runs}}\approx 10^{7}$ (blue line), which, again, is above our measurement budget. Nevertheless, we decrease the regularization parameter $\beta$, which leads to larger error bars, but also allows to learn some of the spatial disorder in $H_{\mathrm{XY}}$, as is shown in Fig. 8b for $\beta=2$ and $N_{\mathrm{runs}}=1.7\times 10^{6}$ (orange data). Note that already before reaching a plateau for $\beta=2$ in Fig. 8a (orange line) the error bars of $\bm{c}^{\mathrm{rec}}$ become very small, which can be seen in Fig. 8b for $\beta=2$ and $N_{\mathrm{runs}}=6.6\times 10^{6}$ (orange data). This suggest to further reduce $\beta$. Ultimately, this process of subsequently reducing $\beta$ and increasing $N_{\mathrm{runs}}$ allows us to learn $H_{\mathrm{XY}}$ up to statistical errors, and leads to $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}\sim N_{\mathrm{runs}}^{-1/2}$ for the entire range of $N_{\mathrm{runs}}$.
Figure 9: Learning collective dissipation using the additional constraints defined in the main text, and regularization strength $\beta=2$. (a) The error of the additional constraints $\Delta^{\mathrm{add}}$ as introduced in Eq. (52) for different ansätze for the dissipation as described the main text. (b) Snapshots of the learned dissipation rates $d_{k}^{\mathrm{rec}}$ as indicated by the asterisks in the left panel. Black crosses indicate the true dissipation rates. All error bars are computed via Bootstrapping as explained in Appendix D, using $10$ samples. Integrals for dissipation correction are evaluated as for Fig. 8. We further restrict $d_{-},d_{z},d_{z}^{\mathrm{col}}\geq 0$ and numerically search for the optimal rates within the bounds.
$[M_{D}(\bm{d})]_{\kappa,(i,j)}=\frac{1}{2}\sum_{k,l}d_{kl}[M^{(k,l)}]_{\kappa,(i,j)}\\ =\frac{1}{2}\int_{0}^{T}\sum_{k,l}d_{kl}\expectationvalue*{\sigma_{k}^{z}[h_{ij},\sigma_{l}^{z}]+\mathrm{h.c.}}_{\kappa,t}\mathrm{d}t,$ (50)
F.1 Learning collective dissipation